At 9 am a freight train starts from L.A. and heads for Chicago at 50 mph. 2 hours later, a passenger train leaves the same station for chicago traveling 60 mph.
A. What time does the passenger train overtake the freight train?
B. How far are they from L.A. when this happens?

Answer by Paul(988)

Let the time be x
50(x+2)=60(x)
50x+100=60x
10x=100
x=10
10

Hence, it takes 10 hours for it to overtake the frieght tarin.
The tatal distance is 600. [50(10+2)=60(10)], = 600=600